# Quantifying and Mitigating Unimodal Biases in Multimodal Large Language Models: A Causal Perspective

### v1

@@ 2 A Causal Framework
Figure 2: Causal graph of MLLM’s Prediction on VQA problems. We highlight the difference between a cognitively-inspired correct reasoning path ($G_{h}$) and the undesired effects that some factors might have on the model’s prediction (e.g., vision and language biases). By performing controlled interventions of the images ($I$) and of the question ($Q$, $T$), we are able to quantify the causal effects of each factor.
In this section, we first introduce the causal graph of MLLM’s prediction on VQA problems. Then, we leverage the causal graph to elucidate the biases inherent in VQA, especially vision and language biases. Finally, we assess the causal effects of such biases on MLLMs’ prediction by performing controlled interventions Pearl (1995).
@@ 2.2 Causal Graph of MLLMs’ Prediction
${s}=f_{\text{abstract }}({q}),\ \ g=f_{{s}}({i}),$ (1)
where they first abstract the core semantic meaning $s$ from the question $q$ by some cognitive process $f_{\text{abstract }}$, and then relate $s$ to the image $i$ to yield the final result $g$. We present these mechanisms in the green subgraph $G_{h}$ of Figure 2.
In contrast, the causal mechanisms of how a model might solve the same VQA problem $m$ are as follows:
$a=f_{\text{blackBox }}({i},{q}),\vspace{-5pt}$ (2)
where it is uncertain which aspect(s) of $q$ the model considers and how it interacts with the image $i$.
@@
Intuitively, if the model learns the correct causal mechanisms to solve a VQA problem (i.e., the green subgraph $G_{h}$ of Figure 2), it should understand how both the image and the question contribute to the ground-truth result ${G}$ (via ${I}\rightarrow G$ and ${S}\rightarrow G$). Consequently, the model’s predictions should exhibit sensitivity and robustness to changes in the ground truth, namely ${G}\rightarrow A$. No spurious correlations should directly affect $A$ without going through the mediator $G$. Based on the above analyses, we articulate the concept of model sensitivity and robustness on VQA problems:
@@ 2.3 Causal Analysis of VQA Biases
After formulating the cognitively-inspired subgraph $\mathcal{G}_{h}$ and defining the path of unimodal biases, we are able to quantify the causal effects of each causal factor on another by performing controlled interventions Pearl (1995).
@@
    Partially controllable interventions on $Q$. The question $Q$ can be modified in two ways: (i) both $S$ and $T$ are altered, or (ii) $T$ is altered but $S$ remains unaffected.
@@
Next, we explain how we can obtain the causal effect from the interventions. Let us consider an intervention $\operatorname{do}\left(X:x\rightarrow x^{\prime}\right)$, where $X\in\{I,Q,T\}$ and the VQA problem $M=\{I,Q\}$. We denote the distribution before the intervention $\mathbb{P}(A\mid{I},{Q})$ as $P$ and the distribution after the intervention as $P^{\prime}$.
Following the distributional definition of causal effect by Pearl (1995), we quantify the effect of factor $X$ in our causal graph using a distance metric $\delta$ between the distributions $P$ and $P^{\prime}$. That is,
$\operatorname{CE}=\delta\left(P,P^{\prime}\right),$ (3)
where $\operatorname{CE}$ denotes the causal effect and can be further divided into the total causal effect ($\operatorname{TCE}$, i.e., the joint effect through all the directed causal paths from a variable to another), or the direct causal effect ($\operatorname{DCE}$, i.e., the effect of the directed causal path from a variable to another without any intermediate variables) Pearl (2022).
@@
When intervening on the images $I$, we can obtain the size of causal effect of $I$ on $A$, namely
@@
  • The path ${I}\rightarrow G\rightarrow A$ represents the desired decision-making route we aim for the model to adopt, where it responds to alterations in the ground-truth answer.
  • The path ${I}\rightarrow A$ illustrates a spurious correlation that the model might have learned, where it depends on certain visual contexts potentially linked to their prevalence in the training corpus.
We can quantify the $\operatorname{DCE}$ of $I$ on $A$, namely the strength of the path ${I}\rightarrow A$, by maintaining $G$ constant during each intervention on $I$:
@@
As for questions, by intervening on $Q$, we can compute the total causal effect of $Q$ on $A$:
@@
Controlling for the core semantic meaning $S$ will instead allow us to obtain the $\operatorname{DCE}$ of the textual surface form $T$ on $A$:
@@
Note that, since there is no mediator between $T$ and $A$, the $\operatorname{DCE}({T}\rightarrow A$) is also $\operatorname{TCE}$ of $T$ on $A$. Due to that it is generally infeasible to enumerate all possible perturbations of $T$, we could achieve practical results over a certain subset of $T$ by intervening on $Q$ without affecting $S$ Stolfo et al. (2022). Besides, in the context of the VQA problem, we cannot only intervene on $S$ without affecting the textual surface $T$. Nevertheless, the causal influence of $S$ on $A$ can be understood by comparing two quantities we already know, i.e., $\operatorname{TCE}({Q}\text{ on }A)$ and $\operatorname{DCE}({T}\rightarrow A)$.

### v2

@@ 2 A Causal Framework
Figure 2: Causal graph of MLLM’s Prediction on VQA problems. We use the green subgraph $G_{h}$ to represent the desired causal mechanisms and compare it with the undesired effects of vision and language biases. We quantify the causal effects of each factor by performing controlled interventions of the images ($I$) and of the question ($Q$, $T$).
In this section, inspired by Stolfo et al. (2022), we first introduce the causal graph of MLLM’s prediction on VQA problems. Then, we leverage the causal graph to elucidate the biases inherent in VQA, especially vision and language biases. Finally, we assess the causal effects of such biases on MLLMs’ prediction by performing controlled interventions Pearl (1995).
@@ 2.2 Causal Graph of MLLMs’ Prediction
${s}=f_{\text{c}}({q}),\ \ g=f_{{s}}({i}),$ (1)
where the cognitive process $f_{\text{c}}$ is employed to decipher the core semantic meaning $s$ of the question $q$. Then, the function $f_{s}$ is utilized to relate $s$ to the image $i$, yielding the final result $g$. These mechanisms are depicted within the green subgraph $G_{h}$ of Figure 2.
Contrarily, the process of a model to tackle the same VQA problem $m$ involves the following causal mechanisms:
$a=f_{\text{b}}({i},{q}),\vspace{-5pt}$ (2)
where $f_{\text{b}}$ functions as a black box, leaving uncertainty regarding which aspects of $q$ are considered by the model and its interaction with the image $i$.
@@
The essence of correct reasoning lies in the model’s grasp of the underlying causal mechanisms necessary for solving a VQA problem. As illustrated in Figure 2), the green subgraph $G_{h}$ embodies this comprehension. It implies an understanding of how both the image and the question jointly influence the ground-truth result ${G}$ (via ${I}\rightarrow G$ and ${S}\rightarrow G$). Consequently, the model’s predictions should exhibit sensitivity and robustness to changes in the ground truth Stolfo et al. (2022), namely ${G}\rightarrow A$. No spurious correlations should directly affect $A$ unless it goes through the mediator $G$. Based on the above analyses, we articulate the concept of model sensitivity and robustness on VQA problems:
@@ 2.3 Causal Analysis of VQA Biases
Once the desired causal mechanism is formulated and the path of unimodal biases is defined, we are able to quantify the causal effects of each causal factor on another by performing controlled interventions Pearl (1995).
@@
    Partially controllable interventions on $Q$. The question $Q$ can be modified in two ways: (i) altering both $S$ and $T$ , or (ii) altering $T$ but keeping $S$ unaffected.
@@
Next, we explain how we can obtain the causal effect from the interventions. Consider an intervention formulated as $\operatorname{do}\left(X:x\rightarrow x^{\prime}\right)$, where $X\in\{I,Q,T\}$ in the context of the VQA problem defined by $M=\{I,Q\}$. For simplicity, the probability distribution prior to the intervention, $\mathbb{P}(A\mid{I},{Q})$, is denoted by $P$, while the distribution following the intervention is represented as $P^{\prime}$. Following Pearl (1995), we measure the causal effect of each factor in our causal graph by employing the distance metric $\delta$ to gauge the divergence between distributions $P$ and $P^{\prime}$:
$\operatorname{CE}=\delta\left(P,P^{\prime}\right).$ (3)
It is important to note that $\operatorname{CE}$ denotes the causal effect, which can refer to either of the two components: the total causal effect ($\operatorname{TCE}$), signifying the joint effect across all causal paths from one variable to another, and the direct causal effect ($\operatorname{DCE}$), indicating the effect of the directed causal path devoid of intermediary variables Pearl (2022).
@@
The causal effect size of the image on the answer can be obtained by conducting intervention on $I$, i.e.,
@@
  • The desired decision-making route we aim for the model to adopt ( the ${I}\rightarrow G\rightarrow A$ path), where the model responds accordingly to alterations in the ground-truth answer.
  • The spurious correlation that the model might have learned (the ${I}\rightarrow A$ path), where the model depends on certain visual contexts potentially linked to their prevalence in the training corpus.
Maintaining $G$ constant during each intervention on $I$ allows us to quantify the $\operatorname{DCE}$ of $I$ on $A$, which refers to the strength of the path ${I}\rightarrow A$:
@@
As for the total causal effect of the question on the answer, we can quantify it by intervening on $Q$:
@@
Instead of changing the core semantic meaning, maintaining $S$ constant will enable us to derive the $\operatorname{DCE}$ of the textual surface form $T$ on $A$:
@@
It should be noted that the $\operatorname{DCE}({T}\rightarrow A$) equates to the $\operatorname{TCE}$ of $T$ on $A$, due to there is no mediator between $T$ and $A$. Given the impracticality of enumerating all possible perturbations of $T$, practical results within a specific subset of $T$ could be attained by intervening on $Q$ while leaving $S$ unaffected Stolfo et al. (2022). Besides, in the context of the VQA problem, solely intervening on $S$ without influencing the textual surface $T$ is unattainable. However, comprehending the causal impact of $S$ on $A$ becomes feasible through the comparison of two known quantities: $\operatorname{TCE}({Q}\text{ on }A)$ and $\operatorname{DCE}({T}\rightarrow A)$.